OpenStudy (anonymous):

You toss two fair dice, one red and one blue. How many different ways can they land? What is the probability that the sum of the dots on the face of the die is 7?

OpenStudy (xapproachesinfinity):

one die can land in 6 different way so two dice will have 6+6 ways which 12 ways makes sense?

OpenStudy (anonymous):

Yeah I'm following! @xapproachesinfinity

OpenStudy (xapproachesinfinity):

so part one is done

OpenStudy (lochana):

so first we need to find all ways of landing of two dices. all situations = 6^2 = 36 now you need situations where sum of dots is 7 it can be 1+6 2+5 3+3 4+2 5+2 6+1 total = 6 probability = 6/36 =1/6

OpenStudy (lochana):

is it 12? I am confused @xapproachesinfinity

OpenStudy (xapproachesinfinity):

now we are looking for possible pairs that sum up to 7 we have 1+6 6+1 3+4 4+3 2+5 5+2 that makes 6 ways we can have a sum of 7

OpenStudy (xapproachesinfinity):

so portability then is 6/12 chances =1/2

OpenStudy (anonymous):

oh my!

OpenStudy (lochana):



OpenStudy (anonymous):

and i was gong to retire for the evening

OpenStudy (xapproachesinfinity):

oh pellet should be 36 possible outcomes

OpenStudy (xapproachesinfinity):

my bad

OpenStudy (anonymous):

ok whew gn

OpenStudy (lochana):

this is how faces of dices can be turned .@xapproachesinfinity am i wrong?

OpenStudy (xapproachesinfinity):

no i was wrong bad habit of making errors

OpenStudy (xapproachesinfinity):

we 6 by 6 possibilities giving us 36 different outcomes

OpenStudy (anonymous):

Okay! so then the 6/36 would be 1/6 for the probability of the second half?

OpenStudy (xapproachesinfinity):

yes correct

OpenStudy (xapproachesinfinity):

sorry for the first error

OpenStudy (anonymous):

@xapproachesinfinity You're fine! I appreciate your help. I had no clue how to work with the dice so that helped a lot. Thank you!!!

OpenStudy (xapproachesinfinity):

you might be saying why it is 36 if you try to pull all different possible pairs you going to get 36 of them

OpenStudy (anonymous):

Okay, gotcha!

OpenStudy (xapproachesinfinity):

here i want to put the table but cutting time and use the internet is better http://i.stack.imgur.com/9CoBC.gif

OpenStudy (xapproachesinfinity):

this table shows all possible pairings we can have with to dice

OpenStudy (xapproachesinfinity):

somehow the ones that add up to 7 lie in diagonal lol

OpenStudy (xapproachesinfinity):

can you see that?

OpenStudy (xapproachesinfinity):

i guess you couldn't see it lol

OpenStudy (anonymous):

I'm so sorry haha I just saw this! But thank you so much. That makes so much sense!! @xapproachesinfinity